OpenStudy (anonymous):

5. Rationalize the denominator 3/3-2 sqrt 2 +4+ sqrt 2

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{3-2\sqrt{2}+4\sqrt{2}}\quad ?\)

OpenStudy (jdoe0001):

shoot.. ahemm

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{3-2\sqrt{2}+4+\sqrt{2}}\quad ?\)